# ShareChat: A Dataset of Chatbot Conversations in the Wild

### v1

@@ Abstract
While Large Language Models (LLMs) have evolved into distinct platforms with unique interface designs and capabilities, existing public datasets treat models as generic text generators, stripping away the interface context that actively shapes user interaction. To address this limitation, we present ShareChat, a large-scale, cross-platform corpus comprising 142,808 conversations and over 660,000 turns collected from publicly shared URLs across five major platforms: ChatGPT, Claude, Gemini, Perplexity, and Grok. ShareChat distinguishes itself by preserving native platform affordances often lost in standard logs, including reasoning traces, source links, and code artifacts, while spanning 101 languages over the period from April 2023 to October 2025. Furthermore, ShareChat offers substantially longer context windows and greater interaction depth than prior datasets. We demonstrate the dataset’s multifaceted utility through three demonstrative use cases: (1) analyzing conversation completeness to measure user intent satisfaction; (2) evaluating source citation behaviors in content generation; and (3) conducting temporal analysis to track evolving usage patterns. This work provides the community with a vital and timely resource for understanding authentic user-LLM chatbot interactions in the wild
@@
1 Introduction
Conversational large language model (LLM)-based chatbot services have evolved rapidly in the past three years. The first widely adopted general purpose LLM chatbot, ChatGPT, was launched in November 2022 and has reached more than 700 million weekly active users by mid 2025 OpenAI Research (2025). Following this success, many companies quickly released their own chatbot services, which also attracted substantial public attention. For example, Anthropic introduced the Claude families in 2023 Anthropic (2023, 2024, 2025), Google deployed the Gemini family of models Anil et al. (2023), xAI launched Grok as a chatbot integrated with the social media platform X xAI (2023), Perplexity emerged as an answer engine that combines conversational interaction with web search Perplexity AI (2024), and DeepSeek released its R1 reasoning model and corresponding consumer chatbot app in January 2025 DeepSeek AI (2025). Although all of these services are built on text-based LLMs, they differ in interface design, supported features, and safety policies, which in turn shape how users interact with them.
For example, Grok can surface live posts from X when providing answers, Claude models are optimized and evaluated for coding, math, and analysis tasks Anthropic (2023, 2024), and Perplexity consistently presents responses with explicit source citations Perplexity AI (2024). These platform level design choices are not static. Commercial systems typically undergo continuous reinforcement learning and fine tuning on conversation logs and human feedback Ouyang et al. (2022); Bai et al. (2022); Chen et al. (2024), which tends to reinforce platform specific strengths and norms over time. As a result, it is important to understand interaction dynamics and user experience across platforms, instead of treating LLM capabilities as a property of models in isolation.
However, most widely used LLM benchmarks focus on curated single turn tasks, such as multiple choice knowledge tests or coding and math word problems Hendrycks et al. (2021); Cobbe et al. (2021). In practice, the use of conversational LLMs is much more diverse and complex: users ask follow up questions, change goals, or co construct content over many turns. Recent work has shown that even state of the art models often get “lost” in multi turn conversations, with reliability degrading as instructions are refined across turns Laban et al. (2025a). Studying how people actually interact with LLMs in the wild therefore requires large scale, real conversation datasets, rather than only synthetic prompts or short, benchmark style dialogues.
Several real world chatbot LLM conversation datasets have been released in the past two years Zhao et al. (2024); Zheng et al. (2024); Köpf et al. (2023). For instance, WildChat Zhao et al. (2024) compiles about one million conversations between users and ChatGPT collected via an access gateway that offered free usage in exchange for consent to share logs. LMSYS-Chat-1M Zheng et al. (2024) contains one million conversations with twenty five state of the art models recorded through a single web interface (the Vicuna demo and Chatbot Arena). OpenAssistant Conversations Köpf et al. (2023) provides human generated assistant style dialogues and detailed preference annotations to support alignment research. These datasets are important resources, but they have several limitations:
Single-Interface Mediation: They are all mediated via a central interface even when multiple backend models are available. Consequently, they do not fully capture how different commercial platforms, with distinct interaction designs and product positioning, influence the content that users share and the way they engage with LLMs.
Text-Only Representation: In addition, existing real conversation datasets primarily focus on plain text message content. Yet real interactions with contemporary LLM systems can include more structure: users may enable a thinking mode that exposes intermediate reasoning traces, and some platforms attach source links or other metadata to model outputs, such as the cited web pages surfaced by Perplexity or embedded X posts shown by Grok. A richer multi platform dataset should preserve these elements when they are available, so that researchers can better understand how users interpret model reasoning, how they rely on external evidence, and how platform affordances shape behavior across turns.
Limited Interaction Depth: Moreover, although some datasets explicitly include multi turn conversations Zhao et al. (2024), the average depth of a conversation is still relatively short. Given that context length is central to LLM training and performance, longer real user chatbot conversations are valuable both for understanding how user intents and prompts evolve and for investigating when and how models succeed or fail at using extended context Laban et al. (2025a).
Observer Bias: Finally, data collection in these studies typically relies on pre-interaction consent, where users agree to share their data before or during the interaction in exchange for service access. This introduces the Observer Bias (or Hawthorne Effect) McCambridge et al. (2014), as users who are aware they are being monitored for research purposes may consciously or unconsciously alter their behavior, which often exhibits higher social desirability compared to authentic, unobserved private usage. In contrast, ShareChat relies on post-hoc sharing, where the decision to publish is made only after the natural interaction has concluded, thereby preserving the integrity of the original behavior during the conversation itself.
In this study, we address these gaps by constructing a large scale, multi platform corpus of authentic user-LLM conversations. We compile 142,808 conversations with more than 660,293 interaction turns from five widely used platforms: Claude, Perplexity, Grok, ChatGPT, and Gemini. All conversations are collected from URLs that users chose to share publicly on the respective platforms, spanning more than twenty months of AI chat usage (April 2023 to October 2025) and covering 101 distinct languages. The dataset therefore captures authentic usage patterns from a diverse, global user base, rather than simulated prompts or tasks designed by researchers. Each conversation record contains the complete sequence of user and assistant turns, along with platform specific metadata such as, when available, timestamps, model thinking traces, or source links attached by the interface. The data collection is done under IRB approval (#28569).
To demonstrate the specific utility of ShareChat, besides describing the basic statistics of the corpus, we present three illustrative case studies, each designed to leverage ShareChat’s unique features to overcome the limitation of prior resources. First, we perform conversation completeness analysis, where we label whether specific user intentions are fully addressed. This utilizes ShareChat’s high interaction depth to capture the resolution of complex goals often obscured in shorter, shallow exchanges. Second, we carry out source domain analysis that connects assistant responses to external resources surfaced by platforms like Perplexity and Grok. This leverages our preservation of non-textual metadata to evaluate how models ground their claims, which is a dimension absent in text-only datasets. Third, we conduct temporal analysis to examine response time and engagement patterns. This exploits our native turn-level timestamps to reveal authentic real-world interaction rhythms, free from the temporal distortions often introduced by observer bias in monitored settings. Together, these analyses highlight the value of a multi-platform, multi-turn, and multi-lingual conversation corpus for understanding real user experiences with LLMs.
Our contributions are twofold. First, we construct and analyze what is, to our knowledge, the first multi platform, multi turn, and multi lingual conversation dataset of user LLM interactions, with substantially longer conversations on average than prior corpora and lower levels of toxic content. Second, we provide three concrete use cases: conversation completeness analysis, temporal analysis, and source link analysis, which together demonstrate how the dataset can be used to study user interaction behavior in depth.
2 Data Collection
For the five differnet platfomrs, we developed systematic, platform specific data extraction pipelines to collect conversation content, metadata, and platform specific features from publicly accessible shared conversation URLs found on the Internet. Although we found that such URLs are often shared on social media platforms such as X, Reddit p(formerly Twitter), Reddit, and Discord, exhaustively monitoring these individual ecosystems presents significant scalability challenges. To tackle this, we utilized Internet archival services, specifically the Internet Archive (Wayback Machine), to search for and extract URLs matching specific patterns (e.g., “chatgpt.com/share/*” for ChatGPT), This archival-based discovery method is a well-established strategy for aggregating public digital resources, previously employed to compile datasets for web-scale text Raffel et al. (2020), historical news archives Hamborg et al. (2017), and corporate disclosure analysis Boulland et al. (2024). Then, for each platform, data collection covered a distinct time window, as summarized in Table 1.
| Platform | Share URL Format | Collection Period |
| ChatGPT | chatgpt.com/share/* | May 2023 to Aug 2025 |
| Perplexity | perplexity.ai/search/* | Apr 2023 to Oct 2025 |
| Grok | grok.com/share/* | Dec 2024 to Oct 2025 |
| Gemini | gemini.google.com/share/* | Apr 2024 to Sep 2025 |
| Claude | claude.ai/share/* | Not available |
For every shared conversation page, we used automated browser control with Selenium to interact with the page, retrieve the rendered HTML, and parse it into a structured JSON representation. The resulting records include the ordered sequence of conversation turns, user prompts, model responses, and platform specific metadata. When certain elements were only visible after user interaction, we scripted the corresponding actions. For example, to collect the thinking content of Claude, we first locate the thinking bar element and then trigger a click on it when the content is only accessible through interaction.
Filtering of Identifiable Information.
We prioritized user privacy by anonymizing personally identifiable information (PII). We used Microsoft’s Presidio111https://microsoft.github.io/presidio/ as the framework, Spacy222https://spacy.io/ for Named Entity Recognition, and custom rules to identify and remove PII across various data types—such as names, phone numbers, emails, credit cards, and URLs—in multiple languages including English, Chinese, Russian, French, Spanish, German, Portuguese, Italian, Japanese, and Korean.
@@ 2 Data Collection
3 Preliminary Analysis
The final dataset contains conversations from five platforms, featuring a wide range of languages, diverse user prompts and longer conversation with less toxic content compared to other datasets. In this section, we describe the basic statistics of the dataset and platform-specific attributes. We also present the results of our toxicity analysis and topic distribution.
3.1 Basic Statistics
In total, ShareChat has 142,808 conversations with more than 660,293 interaction turns. Table 2 presents comprehensive statistics comparing our multi-platform dataset with existing conversation corpora on number of turns, token length and languages used. The language detection was performed at message-level using lingua-py and then combined into conversation-level statistics in Table 2. Token counts were computed using the Llama-2 tokenizer for consistent cross-platform comparison Zhao et al. (2024).
In summary, ShareChat significantly exceeds existing corpora in conversation depth, averaging 4.62 turns compared to the 2.02 turns observed in LMSYS-Chat-1M. Similarly, it demonstrates greater linguistic diversity by spanning 101 distinct languages, whereas the most diverse competitor covers only 65. The model responses are also substantially longer, with a mean of 1115.30 tokens compared to 215.71 for LMSYS-Chat-1M, suggesting the presence of richer and more substantive interactions. ChatGPT comprises the largest portion of the corpus with 102,740 conversations, offering the broadest language coverage.
The key distinguishing characteristic of ShareChat is the substantially longer and more heterogeneous distribution of conversational turns compared to existing public corpora as shown in Figure 1. Whereas prominent benchmarks such as Alpaca, Dolly, LMSYS-Chat-1M, and WildChat exhibit median turn lengths of 1.0, our aggregated dataset achieves a median of 2.0 turns, indicating that conversations in our corpus are fundamentally more multi-turn in nature. This pattern is consistently maintained across individual platforms with results shown in Appendix A.1.
For language detection, we first identified the language of each message using the lingua library, and we include these per message labels in the final dataset. We then assigned a dominant language to each conversation as the most frequently detected language across the user turns. As shown in Figure 2, English accounts for the clear majority of conversations across all platforms, with Japanese as the second most frequent language and all other languages each contributing a much smaller share. Compared to previous datasets, ShareChat has a more balanced language distribution.
As ShareChat features significantly longer conversation turns and higher token counts per message, it addresses a critical gap in current model development pipelines, which often rely heavily on single-turn instruction tuning. By capturing extended interactions, the data supports training for long-context dependencies and exploring complex user behaviors that only emerge during deep, multi-turn exchanges.
3.2 Platform-Specific Interaction and Metadata Attributes
Different platforms reflect distinct design philosophies. ChatGPT and Gemini offer general-purpose conversational assistance. Perplexity focuses on search-oriented information retrieval. Grok targets social media-integrated question answering. Claude specializes in technical and reasoning support. These structural differences align with intended use cases and shape the types of user interactions preserved in ShareChat. Table 3 outlines the distinct data attributes captured across the five platforms, which vary significantly based on their functional priorities.
While ChatGPT and Gemini follow standard conversational formats, other platforms expose specialized content structures. Claude is unique in preserving “thinking blocks” and versioned code artifacts, enabling the analysis of intermediate reasoning and iterative coding processes. In contrast, Perplexity and Grok structure responses around retrieval. Perplexity organizes content into answers, sources, and images with inline citations, while Grok features a dedicated pane for web and social media sources (X posts). These structural differences allow for distinct lines of inquiry: technical workflow analysis for Claude versus citation accuracy and source integration studies for Perplexity and Grok.
@@ 3.2 Platform-Specific Interaction and Metadata Attributes
| Basic Turn Structure | ✓ | ✓ | ✓ | ✓ | ✓ |
| Metadata Capture | ✓ | ✓ | ✓ | ✓ | ✓ |
| HTML/MHTML | ✓ | ✓ | ✓ | - | ✓ |
| Code Artifacts | ✓ | - | - | - | - |
| Thinking Blocks | ✓ | - | ✓ | - | - |
| Analysis Blocks | ✓ | - | - | - | - |
@@
| Image Metadata | - | ✓ | - | - | - |
| Timestamps / turn | - | - | ✓ | ✓ | - |
| User Identifiers | ✓ | ✓ | - | ✓ | ✓ |
| View Counts | - | ✓ | - | - | - |
@@ 3.3 Toxicity Analysis
To assess toxicity in ShareChat, we employed both DetoxifyHanu and Unitary team (2020) and OpenAI Moderation API333https://platform.openai.com/docs/guides/moderation to detect toxic content in the conversations, following the approach used in WildChat analysis Zhao et al. (2024). Detoxify is a pre-trained multilingual toxicity classification model that computes toxicity scores across seven dimensions, while OpenAI Moderation provides commercial-grade content filtering. Given the language coverage limitations of Detoxify, we retained only conversational turns in languages explicitly supported by the model, resulting in a filtered dataset of 949,487 conversations (72.9% of the original corpus) across all five platforms, with more details shown in Appendix A.2. For Detoxify, we applied the same classification threshold as WildChat, marking messages as toxic when the toxicity score exceeds 0.1, while OpenAI Moderation uses its internal flagging mechanism to identify toxic content.
The results for turn-level statistics are shown in Figure 3. Our analysis revealed substantial differences between the two detection methods, consistent with limited agreement observed in prior work where WildChat Zhao et al. (2024) reported that Detoxify flagged 8.12% of user turns compared to the 6.05% flagged by OpenAI Moderation. At the turn level, we observed distinct patterns where Detoxify consistently detected higher toxicity rates for user messages, ranging from 2.8% to 5.6%, compared to the 1.3% to 4.0% range identified by OpenAI Moderation across all platforms. Notably, Claude showed the highest user toxicity under Detoxify detection at 5.6%. There appears to be a positive correlation between user and LLM toxicity rates within platforms. For example, Claude exhibited the highest user toxicity with 5.6% via Detoxify and 2.7% via OpenAI, and similarly showed elevated LLM toxicity at 6.4% and 3.1% respectively. Conversely, Perplexity demonstrated the lowest rates for both users, at 2.8% via Detoxify and 1.3% via OpenAI, and LLMs, at 0.5% via Detoxify and 1.8% via OpenAI.
Notably, ShareChat contains substantially less toxic content compared to existing benchmarks. Our overall user toxicity rates of 4.1% via Detoxify and 2.9% via OpenAI are considerably lower than both WildChat at 6.05% and LMSYS-Chat-1M at 3.08% using the OpenAI method Zhao et al. (2024). Similarly, our LLM toxicity rates of 1.6% via Detoxify and 3.2% via OpenAI are lower than the 5.18% reported by WildChat and comparable to the 4.12% in LMSYS-Chat-1M.

### v2

@@ Abstract
While academic research typically treats Large Language Models (LLM) as generic text generators, they are distinct commercial products with unique interfaces and capabilities that fundamentally shape user behavior. Current datasets obscure this reality by collecting text-only data through uniform interfaces that fail to capture authentic chatbot usage. To address this limitation, we present ShareChat, a large-scale corpus of 142,808 conversations (660,293 turns) sourced directly from publicly shared URLs on ChatGPT, Perplexity, Grok, Gemini, and Claude. ShareChat distinguishes itself by preserving native platform affordances, such as citations and thinking traces, across a diverse collection covering 101 languages and the period from April 2023 to October 2025. Furthermore, ShareChat offers substantially longer context windows and greater interaction depth than prior datasets. To illustrate the dataset’s breadth, we present three case studies: a completeness analysis of intent satisfaction, a citation study of model grounding, and a temporal analysis of engagement rhythms. This work provides the community with a vital and timely resource for understanding authentic user-LLM chatbot interactions in the wild. The dataset will be publicly available.
@@
1 Introduction
Conversational Large Language Model (LLM)-based chatbot services have evolved rapidly in the past three years. The first widely adopted general-purpose LLM chatbot, ChatGPT, was launched in November 2022 and has reached more than 700 million weekly active users by mid-2025 openai2025usagepatterns. Following this success, many companies quickly released their own chatbot services, which also attracted substantial public attention. For example, Anthropic introduced the Claude families in 2023 anthropic2023claude2card; anthropic2024claude3card; Anthropic_2025, Google deployed the Gemini family of models gemini2023report, xAI launched Grok as a chatbot integrated with the social media platform X xai2023grok, and Perplexity emerged as an answer engine that combines conversational interaction with web search perplexityhelp2024.
Although all of these services are built on text-based LLMs, they differ in interface design, supported features, and safety policies, which in turn shape how users interact with them. For example, Grok can surface live posts from X when providing answers, Claude models are optimized and evaluated for coding, math, and analysis tasks anthropic2023claude2card; anthropic2024claude3card, and Perplexity consistently presents responses with explicit source citations perplexityhelp2024. Commercial systems typically undergo continuous reinforcement learning and fine tuning on conversation logs and human feedback ouyang2022instructgpt; bai2022hh; chen2024chatgpt, which tends to reinforce platform-specific strengths and norms over time.
However, current research often fails to capture the complexity of real-world LLM deployment. While commercial platforms have evolved into complex ecosystems with unique features, the scientific community relies on datasets that homogenize these interactions. As seen in both synthetic xu2024magpie; li2025infinity; ding2023enhancing and real-world datasets zhao2024wildchat; zheng2023lmsys; kopf2023openassistant; zhang2025large; taori2023stanford; conover2023free, current corpora systematically strip away interface context, treating diverse products as identical, generic text boxes. For instance, WildChat zhao2024wildchat compiles about one million conversations between users and ChatGPT collected via an access gateway that offered free usage in exchange for consent to share logs. LMSYS-Chat-1M zheng2023lmsys contains one million conversations with twenty five state of the art models recorded through a single web interface (the Vicuna demo and Chatbot Arena). OpenAssistant Conversations kopf2023openassistant provides human generated assistant style dialogues and detailed preference annotations to support alignment research. These datasets are important resources with several limitations:
1.
Artificial Uniformity: Existing datasets typically collect data via a single, generic interface to standardize the process. This distorts reality. By forcing every model into the same neutral interface, these datasets strip away the unique tools that guide user behavior. Users do not interact with a generic “model” in the real world; they interact with a specific product, and scientific benchmarks should reflect this distinction.
2.
Loss of Non-Textual Context: In addition, while current datasets are largely limited to plain text, authentic LLM interactions rely on rich structural elements, such as visible thinking traces and embedded source citations, that are often stripped from standard logs. This prevents researchers from studying how critical product features actually influence user behavior and prompting strategies.
3.
Limited Interaction Depth: Moreover, although some datasets explicitly include multi-turn conversations zhao2024wildchat, the average depth of a conversation is still relatively short. In practice, the use of conversational LLMs is much more diverse and complex: users ask follow up questions, change goals, or co-construct content over many turns. Recent work has shown that even state of the art models often get “lost” in multi-turn conversations, with reliability degrading as instructions are refined across turns laban2025llmslostmultiturnconversation. Given that context length is central to LLM training and performance, longer real user-LLM conversations, extended in both turn count and token length, are valuable both for understanding how user intents and prompts evolve and for investigating when and how models succeed or fail laban2025llmslostmultiturnconversation.
4.
“Forced” Sharing: Finally, most datasets rely on participants who know they are being watched, introducing the Observer Bias (or Hawthorne Effect) mccambridge2014hawthorne. Users who are aware they are being monitored for research purposes may consciously or unconsciously alter their behavior, which often exhibits higher social desirability compared to authentic, unobserved private usage.
In this study, we address these gaps by constructing a large scale, multi-platform corpus of authentic user-LLM conversations. We compile 142,808 conversations with more than 660,293 interaction turns from five widely used platforms: ChatGPT, Perplexity, Grok, Gemini, and Claude, as shown in Tabel 1. All conversations are collected from URLs that users chose to share publicly on the respective platforms, spanning more than thirty months of AI chat usage (April 2023 to October 2025) and covering 101 distinct languages. As ShareChat relies on post-hoc sharing, it therefore captures authentic usage patterns from a diverse, global user base, rather than simulated prompts or tasks designed by researchers. Each conversation record contains the complete sequence of user and assistant turns, along with platform-specific metadata, when available, timestamps, model thinking traces, or source links attached by the interface. Table 2 outlines the distinct data attributes captured across the five platforms, and the details are shown in Appendix A.2. The data collection is done under IRB approval (#28569).
@@ 1 Introduction
Claude |
| Textual Content | ✓ | ✓ | ✓ | ✓ | ✓ |
@@
| Thinking Blocks | - | - | ✓ | - | ✓ |
| Code Artifacts | - | - | - | - | ✓ |
| Analysis Blocks | - | - | - | - | ✓ |
| Turn Timestamps | ✓ | - | ✓ | - | - |
| Model Version | ✓ | - | ✓ | ✓ | - |
| View/Share Counts | - | ✓ | - | - | - |
To demonstrate the specific utility of ShareChat, besides describing the basic statistics of the corpus, we present three illustrative case studies. First, we perform conversation completeness analysis to assess if user intentions are fully addressed, leveraging ShareChat’s high interaction depth to capture complex goal resolutions. Second, we conduct source domain analysis to evaluate how models ground claims, utilizing preserved non-textual metadata to study citation behaviors. Third, we execute temporal analysis using turn-level timestamps, to reveal authentic interaction rhythms and engagement patterns free from observer bias.
Our contributions are twofold. First, we construct the first large-scale corpus of user-LLM interactions that preserves native affordances of multiple platforms, with substantially longer conversations on average and lower levels of toxic content than prior corpora. Second, we provide three concrete case studies, which together demonstrate how the dataset can be used to study user interaction behavior in depth.
2 Data Collection
For the five different platforms, we develop systematic, platform-specific data extraction pipelines to collect conversation content, metadata, and platform-specific features from publicly accessible shared conversation URLs found on the Internet. Although we find that such URLs are often shared on social media platforms such as X (formerly Twitter), Reddit, and Discord, exhaustively monitoring these individual ecosystems presents significant scalability challenges. To tackle this, we utilize Internet archival services, specifically the Internet Archive (Wayback Machine), to search for and extract URLs matching specific patterns (e.g., “chatgpt.com/share/*” for ChatGPT). This archival-based discovery method is a well-established strategy for aggregating public digital resources, previously employed to compile datasets for web-scale text raffel2020t5, historical news archives hamborg2017news, and corporate disclosure analysis boulland2024websites. For each platform, data collection covers a distinct time window, as summarized in Appendix A.1.
For every shared conversation page, we use automated browser control with Selenium to interact with the page, retrieve the rendered HTML, and parse it into a structured JSON representation. The resulting records include the ordered sequence of conversation turns, user prompts, model responses, and platform-specific metadata. When certain elements were only visible after user interaction, we script the corresponding actions. For example, to collect the thinking content of Claude, we first locate the thinking bar element and then trigger a click on it when the content is only accessible through interaction.
Filtering of Identifiable Information.
We prioritize user privacy by anonymizing personally identifiable information (PII). We use Microsoft’s Presidio111https://microsoft.github.io/presidio/ as the framework, Spacy honnibal2020spacy for Named Entity Recognition (NER), and custom rules to identify and remove PII across various data types, such as names, phone numbers, emails, credit cards, and URLs, in multiple languages including English, Chinese, Russian, French, Spanish, German, Portuguese, Italian, Japanese, and Korean.
3 Preliminary Analysis
The final dataset contains conversations from five platforms, featuring a wide range of languages, diverse user prompts and longer conversations with less toxic content compared to other datasets. In this section, we describe the basic statistics of the dataset, and the results of our toxicity analysis and topic distribution.
3.1 Basic Statistics
In total, ShareChat has 142,808 conversations with more than 660,293 interaction turns. Table 1 compares ShareChat with existing corpora across key metrics: turn count, token length, and language diversity. The language detection was performed at message-level using lingua-py and then combined into conversation-level statistics. Token counts were computed using the Llama-2 tokenizer for consistent cross-platform comparison zhao2024wildchat.
In summary, ShareChat distinguishes itself by extending the scope of analysis along two critical dimensions: interaction depth and content volume. First, conversations are substantially longer, averaging 4.62 turns with a heterogeneous distribution that consistently exceeds existing public corpora (Figure 1 and Appendix A.3). Second, model responses are significantly denser, with a mean of 1,115.30 tokens compared to just 215.71 for LMSYS-Chat-1M, suggesting the presence of richer and more substantive problem-solving.
ShareChat also demonstrates greater linguistic diversity by spanning 101 distinct languages, whereas the most diverse competitor covers only 65. For languages covered, as shown in Figure 2, English accounts for the clear majority of conversations across all platforms, with Japanese as the second most frequent language and all other languages each contributing a much smaller share. Compared to previous datasets, ShareChat has a more balanced language distribution.
@@ 3.2 Toxicity Analysis
To assess toxicity in ShareChat, we employ both Detoxify hanu2020detoxify and OpenAI Moderation API222https://platform.openai.com/docs/guides/moderation to detect toxic content in the conversations, following the approach used in WildChat zhao2024wildchat. Detoxify is a pre-trained multilingual toxicity classification model that computes toxicity scores across seven dimensions, while OpenAI Moderation provides commercial-grade content filtering. Given the language coverage limitations of Detoxify, we retain only conversational turns in languages explicitly supported by the model, with more details shown in Appendix A.4. For Detoxify, we apply the same classification threshold as WildChat, marking messages as toxic when the toxicity score exceeds 0.1, while OpenAI Moderation uses its internal flagging mechanism to identify toxic content.
The results for turn-level toxicity are shown in Figure 3, which reveal substantial differences between the two detection methods, consistent with limited agreement observed in prior work zhao2024wildchat. Furthermore, we find a statistically significant positive rank-order correlation between user and LLM toxicity rates within platforms (Spearman’s for Detoxify; for OpenAI). This suggests a “mirroring” effect where platforms receiving more toxic user inputs tend to generate more toxic outputs. For instance, Claude ranks highest in both user (5.6%) and model toxicity (6.4%) under Detoxify.
Most importantly, ShareChat contains substantially less toxic content compared to existing benchmarks. Our overall user toxicity rates of 4.1% via Detoxify and 2.9% via OpenAI are considerably lower than both WildChat at 6.05% and LMSYS-Chat-1M at 3.08% using the OpenAI method zhao2024wildchat. Similarly, our LLM toxicity rates of 1.6% via Detoxify and 3.2% via OpenAI are lower than the 5.18% reported by WildChat and comparable to the 4.12% in LMSYS-Chat-1M.